Two positive charges of 12×10-10C and 8×10-10C are placed 10cm apart find the work done in bringing the charges 4cm closer


1
Expert's answer
2021-06-28T17:14:05-0400

Potential when the charges are 10cm apart

"q_1=12\times 10^{-10}C\\
q_2=8\times 10^{-10}C\\
d_1=10cm=0.1m\\
V_1=k\frac{q_1q_2}{d_1}\\
k=9\times 10^9Nm^2/C^2\\
V_1=9\times 10^9\times \frac{12\times 10^{-10}\times 8\times 10^{-10}}{0.1}\\
V_1=86.4nJ\\"

Potential when the charge is 4cm apart

"d_2=4cm=0.04m\\
V_2=k\frac{q_1q_2}{d_2}\\
k=9\times 10^9Nm^2/C^2\\
V_1=9\times 10^9\times \frac{12\times 10^{-10}\times 8\times 10^{-10}}{0.04}\\
V_2=216nJ\\"

Work done in bringing the charges 4cm closer is

"W=V_2-V_1\\
W=216-86.4=129.6nJ"